# Ensemble Learning for Large Language Models in Text and Code Generation: A Survey

### v1

@@ Abstract
Generative pretrained transformers (GPTs) are the common large language models (LLMs) used for generating text from natural language inputs. However, the fixed properties of language parameters in individual LLMs can lead to inconsistencies in the generated outputs. This limitation also restricts the models’ ability to represent diverse language patterns due to inherent biases. Moreover, many powerful LLMs are closed-source. This prevents organizations from integrating their data into these systems, raising concerns about data privacy and limiting industry applications. Inspired by the successful application of LLM ensemble models in text generation, recent literature has also investigated their potential in code generation. This article reviews these emerging LLM ensemble approaches. Our goal is to enhance readers’ understanding of existing techniques and encourage further research and practical implementation, aiming to expand the real-world applications of LLM ensemble models in both text and code generation. We categorize these approaches into seven main methods: weight merging, knowledge fusion, mixture of experts, reward ensemble, output ensemble, routing, and cascading. From this list, we focus on four methods and models that show strong performance and potential for broader applications. We analyze their modeling steps, training methods, and output features to provide a clear understanding of their capabilities. Our findings highlight the benefits of LLM ensemble techniques. These include better representation of diversity, improved output quality, and greater flexibility in applications. This information offers valuable insights for selecting models for various real-world tasks involving text and code generation, and potentially applying methods to multimodal LLMs.
@@ 1 Introduction
Many applications in language processing have been framed as generative language modeling tasks. One common task is instruction following, which predicts conversational word sequences from language feature sequences. Many generative tasks in natural and programming language processing aim to generate language features from various types of inputs. They are referred to as text and code generation tasks in this article. These tasks cover a wide range of research areas, including mathematical reasoning, question answering (QA), massive multitask language understanding (MMLU), and code generation. These topics have the common goal of generating language features and differ in the forms of inputs.
The popular LLM for text and code generation is the single GPT [1, 2]. In the context of language generation, the GPT is a generative pretrained transformer model that can represent language features using a sequence of text and position embeddings linked to linguistic characteristics. Language generation approaches using GPT have demonstrated the ability to produce highly intelligible and natural text. [2, 3, 4]. However, the generated responses often exhibit noticeable biases [5] compared to majority-voted answers. Lack of consistency of language models in their responses due to biased training data is one of the main reasons for this sensitivity [6].
For instance, in GPT-based instruction following, fine-tuned, context-dependent GPTs are used to represent language feature embeddings based on conversational texts [2]. The embedding transitions through each layer of the transformer are characterized by attention weights. During training, the parameters of the GPT are estimated based on the cross entropy (CE) and human feedback criterion. At synthesis time, given an input sentence and the trained parameters, the most likely language features are predicted using a meta-generation algorithm [7, 8]. Although this process enhances the robustness of parameter estimation and generation, the single-LLM approach has notable limitations.
@@
Firstly, the fixed properties of language parameters in individual LLMs can lead to inconsistencies in output [9] and limit the model’s ability to express diverse language patterns due to inherent biases [10]. We identify this issue stems from two efficiency assumptions: the efficiency of information retrieval within a fixed context window and the efficiency of implicit knowledge representation without direct access to external knowledge bases. Consequently, the generated language features often depend heavily on the LLM, leading to hallucinated information and degraded text quality [11]. In recent years, various techniques have been proposed to address these model-dependent issues. These include introducing improved language models (e.g., the mixture of experts LSTM [12], herd of LLMs [13], and multimodal LLM [14, 15]), enhancing training criteria (e.g., scalable knowledge distillation training [16]), and modifying the meta-generation algorithm (e.g., increasing computational capacity using intermediate decoding steps [17]). Secondly, many powerful LLMs have closed-source code, which prevents organizations from integrating their data into these systems and raises data privacy concerns, limiting industrial applications.
@@
Since 2022, LLM ensemble learning has emerged as a significant area of machine learning (ML) research, attracting the attention of many language modeling researchers [18, 19]. LLM ensemble learning refers to a class of ML techniques that exploit multiple LLMs of combined information processing to enhance feature generation, pattern recognition and classification. Researchers have explored both architecture-level and model-level ensemble methods. Architecture-level methods include techniques like weight merging [19], knowledge fusion [20], mixture of experts (MoEs) [21], and reward ensembles [22]. Model-level methods encompass output ensemble [23], routing [24], and cascading [9].
Strictly speaking, a MoE is a specific type of transformer model that incorporates multiple expert layers and serves as a base model to be combined in various LLM ensemble models (e.g., output ensemble models). MoEs, as self-attention models, often outperform conventional single LLMs, such as LLaMA 2 70B, while using fewer parameters [21]. Also, LLM ensemble models do not strictly adhere to traditional ensemble learning methods like bagging or boosting, but they function similarly to bias reduction models (e.g., weight merging [19] as bagging [25], cascading [9] as boosting [26]). Given their similar traits to conventional methods, we define our taxonomies as LLM ensemble in this article.
@@
One example of the successful applications of LLM ensembling is code generation. In this approach, an ensemble of LLMs is introduced to replace a single LLM for evaluating the prediction of code snippets [21, 24, 27], as in Figure 1. LLM ensemble techniques have also been applied to the text generation recently to deal with the limitations of the conventional approaches [19, 18, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 20, 44, 45, 46, 22, 47, 48, 49, 50, 51, 23, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 24, 9, 62, 63]. Different with the LLM ensembling in code generation is the data and model evaluation, these emerging language modeling approaches for text generation adopted various model structures. Some of them focus on improving the functions of weight merging [19, 18, 28, 29, 30, 31, 32, 33, 34], knowledge fusing [20, 44], or reward modeling [22, 47, 48, 49] using source LLMs or improving the gating functions to expert layers using MoEs [6, 64]. While some others combine the LLM output responses directly [23, 52, 53, 54, 55, 56] or model the entire routing process from input prompt to LLMs [24, 59, 60, 65, 58, 61, 57, 9, 62].
@@
LLM ensemble techniques outperform single-LLM models, like traditional GPTs, in representing diversity among LLM agents and output features. These techniques effectively combine varied strengths and features to enhance target responses, particularly in tasks like question answering. They utilize architecture-level frameworks, such as MoEs, and model-level frameworks, like output ensemble and routing. Additionally, human developers often use hierarchical structures to optimize from multiple solutions in code production [66], highlighting the need for a review of various LLM ensemble models for code generation applications.
This survey defines and describes LLM ensemble techniques based on original categorizations and subjectively evaluates selected models for their generation performance. From this categorization, we select four methods and models demonstrating strong performance and potential for broader applications. We analyze these models for text and code generation across three aspects: input and target features, model structures and training, and comparisons between text and code generation. This review seeks to improve readers’ understanding of current techniques and encourage further research and practical implementation. We aim to broaden the real-world applications of LLM ensemble models in text and code generation, overcoming the limitations with single-LLM approaches.
While a few surveys touch on related topics, such as the work by Lu et al. [67] on LLM collaboration methods and Yang et al. [68] on ensemble techniques of LLMs and multimodal LLMs across various machine learning fields, our focus is more concentrated. We concentrate specifically on LLM ensemble methods for text and code generation, providing in-depth insights for practical applications.
This article first reviews the conventional and popular transformer framework for language generation, including GPT-based discriminative and generation tasks. We then analyze the limitations of these approaches and introduce key models and techniques relevant to LLM ensemble learning, such as MoEs, output ensemble, and routing. Finally, we review emerging language generation approaches using LLM ensemble techniques, discussing their motivations and implementations, and highlighting remaining challenges and future directions in this field.
@@ 2 Conventional Text Generation Using an LLM
The transformer model, introduced in 2017 [69], marked a significant advancement in natural language generation (NLG). This model is versatile, supporting various generative tasks across time series analysis, vision, and image processing. In NLG, transformers model the relationship between text and its linguistic realizations, enhancing performance across a broad range of applications. For instance, BERT utilizes bidirectional encoder representations to better understand sentence context [70], T5 treats all NLP tasks as text-to-text problems, facilitating its application in translation, summarization, and question answering [71], and Transformer XL extends the ability to learn dependencies beyond fixed lengths, improving handling of longer sequences [72].
Among these, the context-dependent transformers, specifically fine-tuned for particular tasks, have shown the success in NLG [73]. The Generative Pre-trained Transformer (GPT) is a prominent example of such models, known as a large language model. GPT models in language tasks can generate or classify a sequence of text using attention mechanism layers followed by blocks with high-dimensional inner states. The attention mechanism computes a weighted representation of input sequences, focusing on relevant contexts and dependencies to enhance language understanding and generation. Each block contains a feed-forward neural network (FFN) that processes the output from the attention mechanism to learn complex representations. An architectural example of a GPT is illustrated in Figure 2. The architecture of a GPT involves text and position embeddings that represent observed features, which are inputted into the transformer, and the transitions between layers are characterized by attention weights.
Figure 2: Architecture of a GPT as a large language model, which is fine-tuned on a specific task given the text and position embeddings.
@@
The GPT-based language understanding system is typically designed to tackle either discriminative or generative tasks, and its capability to generate human-like responses is limited. To enhance conversational models, multiple approaches are often integrated. One effective method involves applying reinforcement learning from human feedback (RLHF) after initially fine-tuning the model on instruction-following tasks [2]. This process includes using demonstration data consisting of desired response texts provided by humans in response to prompts. A reward model is then trained using the sampled human-ranked outputs. Subsequently, a proximal policy optimization algorithm [78] is used to refine the policy based on the rewards determined by the reward model.
@@ 3 Primary Methods for LLM Ensembling
Since 2022, techniques for ensemble learning with LLMs have been effectively used to model language features. These applications include extracting product features for e-commerce [42], generating image captions [79], and classifying medical diagnoses [80]. One major benefit of ensemble learning is its ability to generalize diverse outputs by combining knowledge through different frameworks. This can be done at the architecture level, such as weight merging and knowledge fusion, or at the model level, like output ensemble and routing (Figure 3). Language generation of LLMs is fundamentally a next-token prediction task. The goal of ensemble language modeling is to improve the quality of predicted tokens. In this section, we review basic methods from the perspective of ensemble techniques.
@@ 3.1.1 Weight Merging
Weight merging refers to the method of combining the weights of multiple trained models into a single model. This method requires architectural modification and can model the diversity among a set of transformer weights using multiple LLMs.
In traditional weight merging, weights from $K$ LLMs are merged where the merge function over the weights $\text{W}_{i}$ for each LLM is typically given by averaging as
@@
where W is the final model weight. Figure 4 shows weight merging using three source LLMs. $K$ can be simply the number of all LLMs [28] or the top-$K$ LLMs can be determined e.g., by greedy algorithm [18]. W can be further interpolated using pre-trained weights and a parameter $\alpha$ [19]. These weight averaging methods, interpolating between models, are shown to improve accuracy on the fine-tuning tasks.
@@
Building on those findings, the extended merging methods extrapolate weights between models [29, 30]. In this framework, a task vector $T$ is typically defined as
@@
where $\text{W}_{pre}$ is the weights of a pre-trained LLM, $\text{W}_{ft}$ is the weights of the same model after fine-tuning, and $T$ is transformed and applied to merge weights from multiple LLMs. Given a set of LLMs, $T$ can be transformed based on the formula $T_{new}$ by various network operation techniques, such as arithmetic operation [29], trim, elect, sign [31], random drop and re-scale [32]. Examples of the arithmetic operations are illustrated in Figure 5. The merge function over the task vector $T_{new}$ can then be calculated as
@@
where $\text{W}_{pre}$ is the weights of a base pre-trained LLM and the scaling term $\lambda$ is determined using held-out validation sets.
@@
Another approach of perplexity-based merging can be found in [33]. This employs a test-time aggregation approach using importance weights to combine the output logits of LLMs during inference; the perplexity over the input prompt was first validated to measure each LLM’s expertise, and then a greedy algorithm was used for perplexity minimization [33] to derive importance weights. This approach extended the weight merging approach to the one at test time. Thus, this training-free ensemble model can easily include newly released open-source LLMs. Because the application of model merging strategies is an expanding area of research, a comprehensive, open-source library has been developed to support them by MergeKit [34].
@@ 3.1.2 Knowledge Fusion
Knowledge fusion is the architecture-level technique to represent the collective knowledge of LLMs by defining different fusing functions and computing the probabilistic distribution matrices of source LLMs [20]. The fusing process of three source LLMs to train a target LLM is shown in Figure 6. This method extends the knowledge distillation technique [81] widely used for text classification, where it transfers knowledge from a larger, more complex model (teacher) to a smaller, simpler model (student) to create an efficient student model.
@@
The problem setting in knowledge fusion is that improve a target LLM by training with a fused matrix $\textbf{P}_{q}$. Given a set of probabilistic distribution matrices $\{\textbf{P}_{q}\}^{K}_{j=1}$ of $K$ LLMs for text sample $q$ from the corpus $C$, $\textbf{P}_{q}$ is obtained by
@@
where $\theta_{j}$ is the parameters of the $j$th LLM and $\textbf{P}^{\theta_{j}}_{q}$ is made by aggregating token-level predictions. Fusion($\cdot$) can be selected among various strategies, e.g., outputs $\textbf{P}_{q}$ using the minimum or weighted average of cross-entropy scores [20]. While training a target LLM using the fusion algorithm, the objective function for transferring the capabilities of source LLMs to the target LLM is derived as
@@
where $\mathbb{D}(\cdot)$ is the discrepancy function between two matrices (e.g., KL divergence), $\textbf{Q}_{q}$ is the output distribution matrix of the target LLM, $\textbf{O}_{q}$ is the one-hot label matrix, and $\lambda$ is the weight parameter.
For LLMs of different size than the target LLM, this knowledge fusion method can be extended to the pairwise method. For each LLM, fixing as a pivot, one can compute pairwise matrices between the pivot LLM and each of the rest source LLMs [44]. The pairwise fused matrix, e.g., $\textbf{P}^{j}_{q}$ can be derived using the pairwise knowledge fusion function as
@@
The pairwise fusion objective for each target LLM is then defined as
@@
where $L_{Fusion}$ is substituted to $L$ in formula (9) to compute the final objective function.
Different LLMs may use different tokenizers, which can result in varied size of their vocabulary matrices. For instance, LLM1’s tokenizer might produce the token ”immediate,” wheres LLM2’s tokenizer generates ”immediately.” This difference in their vocabulary matrices hinders the mathematical fusion process, known as the token alignment problem [82].
To solve this issue, we can use dynamic programming. This technique helps minimize the total cost of editing one sequence of tokens to match another, using a one-to-one exact mapping [82]. Alternatively, we can use a minimum edit distance approach. This method aligns tokens from different tokenizers based on the minimum number of edits needed, which helps maintain important information in the vocabulary distribution [20].
Another way to tackle the token alignment problem is through an embedding representation approach. This method combines collaboration weights with relative representation matrices from the LLMs. These matrices are based on the similarities between tokens, using anchor tokens for reference. By merging these matrices with collaborative model weights, we improve the model’s ability to generalize, making it more effective in tasks like word completion [83, 84].

### v2

@@ Abstract
Generative Pretrained Transformers (GPTs) are foundational Large Language Models (LLMs) for text generation. However, individual LLMs often produce inconsistent outputs and exhibit biases, limiting their representation of diverse language patterns. The closed-source nature of many powerful LLMs further restricts industry applications due to data privacy concerns. Inspired by successes in text generation, LLM ensemble techniques are now increasingly explored for code generation. This article reviews these emerging ensemble approaches to enhance understanding, encourage further research, and promote practical implementation in both text and code generation. We categorize LLM ensembles into seven main methods—weight merging, knowledge fusion, mixture-of-experts, reward ensemble, output ensemble, routing, and cascading—analyzing capabilities of those approaches. Our findings highlight key benefits such as improved diversity representation, enhanced output quality, and greater application flexibility. These insights aid model selection for real-world tasks and crucially, lay groundwork for extending ensemble strategies to multimodal LLMs.
@@ 1 Introduction
Many language processing applications are framed as generative modeling tasks, such as instruction following, which predicts conversational word sequences. This article focuses on text and code generation tasks, which encompass diverse research areas like mathematical reasoning, question answering (QA), massive multitask language understanding (MMLU), and code generation. These fields share the goal of generating language features from varied inputs.
Models based on the Generative Pretrained Transformer (GPT) architecture are widely used for text and code generation [1, 2]. (In this paper, ”GPT” refers to this architectural concept unless a specific model like GPT-4o is named.) These GPT-based LLMs utilize embeddings to represent language features and can generate highly intelligible, natural text [2, 3, 4]. However, their responses often exhibit noticeable biases [5] and inconsistencies, a sensitivity largely attributed to biased training data [6]. This limitation in individual models motivates the exploration of ensemble approaches.
In typical GPT-based instruction following, for example, models are fine-tuned on conversational texts, with attention mechanisms characterizing embedding transitions through layers [2]. Training optimizes parameters via criteria like cross-entropy and human feedback. During synthesis, a meta-generation algorithm predicts likely language features [7, 8]. While this process improves robustness, the inherent limitations of relying on a single LLM persist, underscoring the need for alternative strategies like ensembling.
@@
Individual LLMs face two key limitations. Firstly, their fixed language parameters can cause output inconsistencies [9] and restrict diverse language expression due to inherent biases [10]. We attribute this issue stems from two assumptions about efficient information retrieval within fixed context windows and implicit knowledge representation without external access. Consequently, generated outputs can be model-dependent, leading to hallucinations and degraded text quality [11]. Various techniques address these issues, including improved model architectures (e.g., Mixture-of-Experts [12], model herds [13], multimodal LLMs [14, 15]), enhanced training criteria like knowledge distillation [16], and modified meta-generation algorithms incorporating intermediate decoding steps [17]. Secondly, the closed-source nature of many powerful LLMs hinders data integration and raises privacy concerns, limiting industrial applications.
@@
Since 2022, LLM ensemble learning has rapidly gained prominence in machine learning (ML) research [18, 19]. These techniques combine multiple LLMs to improve performance in tasks such as feature generation, pattern recognition, and classification. Researchers have explored methods at both the architecture-level—such as weight merging [19], knowledge fusion [20], mixture-of-experts (MoE) [21], and reward ensembles [22]—and the model-level, including output ensembles [23], routing [24], and cascading [9].
While MoE is a specific transformer architecture with multiple expert layers—often outperforming larger single LLMs like LLaMA 2 70B with fewer parameters [21]—it can also serve as a base model within broader LLM ensemble frameworks (e.g., output ensembles). LLM ensembles, though not strictly adhering to traditional methods like bagging or boosting, exhibit similar functionalities: weight merging, for instance, resembles bagging in reducing variance [19, 25], while cascading can be seen as a form of boosting [9, 26]. Given these functional parallels, we categorize these approaches as ”LLM ensemble” in this article.
@@
LLM ensembling has shown notable success, for instance, in code generation where multiple LLMs are used to evaluate code snippet predictions more effectively than a single LLM [21, 24, 27] (see Figure 1). Recently, these techniques have also been extensively applied to text generation to address limitations of conventional approaches [19, 18, 21, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 20, 44, 45, 46, 22, 47, 48, 49, 50, 51, 23, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 24, 9, 62, 63]. While differing from code generation applications in data and evaluation specifics, these text generation ensembles utilize diverse model structures. Some focus on enhancing mechanisms such as weight merging [19, 18, 28, 29, 30, 31, 32, 33, 34], knowledge fusion [20, 44], or reward modeling [22, 47, 48, 49] with source LLMs, or improving MoE gating functions [6, 64]. Others directly combine LLM outputs [23, 52, 53, 54, 55, 56] or model the end-to-end routing process from prompt to LLMs [24, 59, 60, 65, 58, 61, 57, 9, 62].
@@
LLM ensemble techniques generally outperform single GPT-based models by better representing diversity among LLM agents and their output features. These methods effectively combine varied strengths to enhance responses, particularly in tasks like question answering, utilizing both architecture-level (e.g., MoEs) and model-level (e.g., output ensemble, routing) frameworks. The practice of human developers using hierarchical structures to optimize code from multiple solutions [66] further underscores the relevance and need for a comprehensive review of LLM ensemble models, especially for code generation applications.
This survey defines and describes LLM ensemble techniques using original categorizations and subjectively evaluates selected models for generation performance. For each technique, we focus on two to three papers that demonstrate representative performance and broader application potential, analyzing them for text and code generation across metrics, model structures/training, and cross-task comparisons. Our review aims to enhance understanding of current techniques, encourage further research and practical implementation, and ultimately broaden the real-world application of LLM ensembles, overcoming single-LLM limitations.
While existing surveys cover related areas—such as Lu et al. [67] on LLM collaboration and Yang et al. [68] on broader ensemble techniques across machine learning fields including multimodal LLMs—our work offers a more concentrated focus. This survey specifically examines LLM ensemble methods for text and code generation, providing in-depth insights relevant to practical applications in these domains.
This article first reviews the conventional transformer framework for language generation, including GPT-based discriminative and generative tasks. We then analyze the limitations of these single-model approaches and introduce key LLM ensemble techniques, such as MoE, output ensembles, and routing. Finally, we survey emerging language generation methods using LLM ensembles, discussing their motivations, implementations, and highlighting remaining challenges and future research directions.
@@ 2 Conventional Text Generation Using an LLM
The transformer model, introduced in 2017 [69], marked a significant advancement in natural language generation (NLG). This model is versatile, supporting various generative tasks across time series analysis, vision, and image processing. In NLG, transformers model the relationship between text and its linguistic realizations, enhancing performance across a broad range of applications. For instance, BERT utilizes bidirectional encoder representations to better understand sentence context [70], Transformer XL extends the ability to learn dependencies beyond fixed lengths [71], XLNet improves upon these approaches with permutation-based training that maintains bidirectionality while avoiding prediction-perception discrepancy [72], Longformer addresses document-level tasks by implementing an attention mechanism that scales linearly with sequence length [73], and T5 treats all NLP tasks as text-to-text problems, facilitating its application in translation, summarization, and question answering [74].
Among these, the context-dependent transformers, specifically fine-tuned for particular tasks, have shown the success in NLG [75]. The Generative Pretrained Transformer (GPT) is a prominent example architecture of such models, known as a large language model. GPT-based models in language tasks can generate or classify a sequence of text using attention mechanism layers followed by blocks with high-dimensional inner states. The attention mechanism computes a weighted representation of input sequences, focusing on relevant contexts and dependencies to enhance language understanding and generation. Each block contains a feed-forward neural network (FFN) that processes the output from the attention mechanism to learn complex representations. An architectural example of a Transformer as a backbone of LLM is illustrated in Figure 2. The GPT architecture involves text and position embeddings that represent observed features, which are inputted into the transformer, and the transitions between layers are characterized by attention weights.
Figure 2: Architecture of a Transformer as a backbone of LLM, which is fine-tuned on a specific task given the text and position embeddings.
@@
The GPT-based language understanding system is typically designed to tackle either discriminative or generative tasks, and its capability to generate human-like responses is limited. To enhance conversational capabilities, multiple approaches are often integrated. Instruction tuning - fine-tuning models on diverse instruction-following tasks - has emerged as a foundational technique [80]. Building on this foundation, reinforcement learning from human feedback (RLHF) provides further refinement [2]. This multi-stage process begins with demonstration data where human experts provide exemplary responses to prompts. Next, a reward model is trained on human-ranked outputs to evaluate response quality. Subsequently, a proximal policy optimization algorithm [81] is used to refine the policy based on the rewards determined by the reward model.
@@ 3 Primary Methods for LLM Ensembling
Since 2022, LLM ensemble learning techniques have been effectively applied to model language features across diverse applications, such as extracting e-commerce product features [42], generating image captions [82], and classifying medical diagnoses [83]. A key benefit of ensembling is its capacity to generalize from diverse outputs by combining knowledge through various frameworks, either at the architecture level (e.g., weight merging, knowledge fusion) or the model level (e.g., output ensemble, routing), as illustrated in Figure 3. Fundamentally, LLM language generation is a next-token prediction task, and ensemble language modeling aims to enhance the quality of these predicted tokens. This section reviews foundational ensemble techniques from this perspective.
@@ 3.1.1 Weight Merging
Weight merging combines the parameters (weights) of multiple trained LLMs into a single model. This architecture-level technique operates directly on the model parameters, typically requiring the source LLMs to share compatible architectures, and aims to consolidate their diverse learned knowledge.
Traditionally, weights $\text{W}i$ from $K$ LLMs are merged, often by averaging:
@@
where W represents the final model’s weights (Figure 4 illustrates this with three source LLMs). $K$ can denote all LLMs [28] or a subset, such as top-$K$ models selected via a greedy algorithm [18]. Furthermore, W can be interpolated with pre-trained weights using a parameter $\alpha$ [19]. These weight averaging methods, by interpolating between models, have been shown to improve accuracy on downstream tasks.
@@
Building on these findings, extended merging methods extrapolate weights by utilizing task vectors [29, 30]. A task vector $T$ is typically defined as the difference between fine-tuned weights ($\text{W}_{ft}$) and pre-trained weights ($\text{W}_{pre}$) of an LLM:
@@
This task vector $T$ can then be transformed (denoted $T_{new}$) using various techniques such as arithmetic operations [29], trimming and sign-based selection (TIES) [31], or random dropping and re-scaling (DARE) [32], as illustrated for arithmetic operations in Figure 5. The final merged weights W are obtained by applying this transformed task vector to a base pre-trained LLM’s weights $\text{W}_{pre}$ with a scaling factor $\lambda$, often determined using a validation set:
@@
Another approach, perplexity-based merging, performs test-time aggregation by combining output logits using importance weights [33]. In this method, each LLM’s expertise is first measured by its perplexity on an input prompt, and a greedy algorithm then minimizes perplexity to derive these importance weights. This extends weight merging to inference time, creating a training-free ensemble that can easily incorporate newly released open-source LLMs. Recognizing the expanding research in model merging, MergeKit offers a comprehensive open-source library to support these strategies [34].
@@ 3.1.2 Knowledge Fusion
Knowledge fusion, an architecture-level technique, aims to consolidate the collective knowledge of multiple source LLMs into a target LLM by defining fusing functions over their probabilistic output distributions [20]. Figure 6 illustrates this process using three source LLMs to train a target LLM. This method extends traditional knowledge distillation [84], which transfers knowledge from a larger teacher model to a smaller student model, primarily for tasks like text classification, to create more efficient student models.
@@
In knowledge fusion, the objective is to enhance a target LLM by training it with a fused probabilistic distribution matrix $\textbf{P}_{q}$. For each text sample $q$ from a corpus $C$, $\textbf{P}_{q}$ is derived by combining the token-level prediction distributions $\textbf{P}^{\theta_{j}}_{q}$ from $K$ source LLMs (where $\theta_{j}$ denotes the parameters of the $j$-th LLM):
@@
The $\text{Fusion}(\cdot)$ function can employ various strategies, such as selecting the output based on the minimum cross-entropy or a weighted average of cross-entropy scores from the source models [20]. During the training of the target LLM, the capabilities of the source LLMs are transferred using the following objective function:
@@
Here, $\mathbb{D}(\cdot)$ is a discrepancy function (e.g., KL divergence) between two matrices, $\textbf{Q}_{q}$ is the output distribution matrix of the target LLM, $\textbf{O}_{q}$ is the one-hot label matrix, and $\lambda$ is a weighting parameter.
This knowledge fusion method can be extended to a pairwise approach, particularly useful when source LLMs differ in size from the target LLM. For each source LLM, designated as a pivot, pairwise fused matrices are computed between this pivot and every other source LLM [44]. The pairwise fused matrix for the $j$-th LLM (when the $v$-th LLM is the pivot), $\textbf{P}^{j}_{q}$, is derived as:
@@
The pairwise fusion objective for the target LLM, incorporating this specific $\textbf{P}^{j}_{q}$, is then defined as:
@@
This $L_{Fusion}$ is subsequently substituted into the overall loss function $L$ in Equation (9) to compute the final training objective.
Different LLMs often employ distinct tokenizers, leading to variations in their vocabulary matrices. For example, one LLM’s tokenizer might output ”cryptocurrency” as a single token, while another’s might produce ”crypto” and ”currency” as separate tokens. This discrepancy in vocabulary representations complicates the mathematical fusion process and is known as the token alignment problem [85].
To address this token alignment issue, dynamic programming can be employed to find a one-to-one exact mapping that minimizes the total cost of transforming one token sequence into another [85]. Alternatively, a minimum edit distance approach aligns tokens from different tokenizers by minimizing the necessary edits, thereby preserving crucial information within the vocabulary distribution [20].
Alternatively, an embedding representation approach tackles token alignment by combining collaboration weights with relative representation matrices from LLMs. These matrices capture token similarities, often anchored by reference tokens, and merging them with collaborative model weights enhances generalization for tasks like word completion [86, 87].